The mass of a proton is 1.67x 10-27 kg (10%) Problem 10: A 50% Part (a) Find the rest energy in joules. 0
Answer #1

Energy , E = mc2 =( 1.67 * 10-27) ( 3*108) 2

= 1.50 * 10-10 joules.

this is the rest energy of protons.
Answer #2

SOLUTION :


Rest mass of proton, m = 1.67*10^(-27) kg 


Speed of light , c = 3*10^8 m/s


So, rest energy of proton 

= m c^2

= rest mass of proton * (speed of light)^2 

= 1.67*10^(-27) * (3*10^8)^2

= 1.503*10^(-10) J  (ANSWER)


Note :

We can convert it into MeV also, as under :

= 1.503*10^(-10) / (1.602*10^(-19))  eV .    [ since 1 eV = 1.602*10^(-19) J ]

= 938.2 * 10^6 eV

= 938.2 MeV

answered by: Tulsiram Garg